OpenStudy (anonymous):

solve for x: 3x - 2x = 65

OpenStudy (amistre64):

add the right side ...

OpenStudy (anonymous):

x=65

OpenStudy (amistre64):

... or left depending on how dyslexic you are :)

OpenStudy (anonymous):

simplify lhs

OpenStudy (anonymous):

the x is raisepower 2

OpenStudy (amistre64):

you should rewrite the equation then, it would help out alot

OpenStudy (anonymous):

which x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[3^{x} - 2^{x} = 65\]

OpenStudy (amistre64):

nice use of the equation editor, but I still think its not quite what you what to present. but I havfe been known to be wrong before :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=3^x%E2%88%922^x%3D65 wolfram says 4 if it is :)

OpenStudy (anonymous):

@amistre that is the question

OpenStudy (anonymous):

\[3^x - 2^x = 65\] \[\ln{(3^x - 2^x)} = \ln{(65)}\] and then?

OpenStudy (anonymous):

thats all